# Hodge diamonds of the Landau–Ginzburg orbifolds

### 26 Feb 2024 v2

@@ 5. Hodge diamond of LG orbifolds
$[x_{1}^{2}x_{2}^{2}x_{3}^{2}x_{4}^{2}x_{5}^{2}]\xi_{\mathrm{id}},[x_{1}^{2}x_{2}^{2}x_{3}^{2}x_{4}^{3}x_{5}]\xi_{\mathrm{id}},[x_{1}^{2}x_{2}^{2}x_{3}^{2}x_{4}x_{5}^{3}]\xi_{\mathrm{id}},[x_{1}^{3}x_{2}^{3}x_{3}^{3}x_{4}]\xi_{\mathrm{id}},[x_{1}^{3}x_{2}^{3}x_{3}^{3}x_{5}]\xi_{\mathrm{id}},$
$[(x_{1}+x_{2}+x_{3})^{3}x_{4}^{3}x_{5}](\xi_{(1,2,3)}+\xi_{(1,3,2)}),[(x_{1}+x_{2}+x_{3})^{3}x_{4}x_{5}^{3}](\xi_{(1,2,3)}+\xi_{(1,3,2)}),[(x_{1}+x_{2}+x_{3})x_{4}^{3}x_{5}^{3}](\xi_{(1,2,3)}+\xi_{(1,3,2)}),$
$[(x_{1}+x_{2}+x_{3})^{3}x_{4}^{2}x_{5}^{2}](\xi_{(1,2,3)}+\xi_{(1,3,2)}),[(x_{1}+x_{2}+x_{3})^{2}x_{4}^{3}x_{5}^{2}](\xi_{(1,2,3)}+\xi_{(1,3,2)}),[(x_{1}+x_{2}+x_{3})^{2}x_{4}^{2}x_{5}^{3}](\xi_{(1,2,3)}+\xi_{(1,3,2)})$
@@
This gives the following diamond
{tikzpicture}

### 25 Mar 2024 v3

@@ 2. $(2)$
$\bigl{[}x_{1}^{2}x_{2}^{2}x_{3}^{2}x_{4}^{2}x_{5}^{2}\bigr{]}\xi_{\mathrm{id}},\qquad\bigl{[}x_{1}^{2}x_{2}^{2}x_{3}^{2}x_{4}^{3}x_{5}\bigr{]}\xi_{\mathrm{id}},\qquad\bigl{[}x_{1}^{2}x_{2}^{2}x_{3}^{2}x_{4}x_{5}^{3}\bigr{]}\xi_{\mathrm{id}},\qquad\bigl{[}x_{1}^{3}x_{2}^{3}x_{3}^{3}x_{4}\bigr{]}\xi_{\mathrm{id}},$
$\bigl{[}x_{1}^{3}x_{2}^{3}x_{3}^{3}x_{5}\bigr{]}\xi_{\mathrm{id}},\qquad\bigl{[}(x_{1}+x_{2}+x_{3})^{3}x_{4}^{3}x_{5}\bigr{]}(\xi_{(1,2,3)}+\xi_{(1,3,2)}),$
$\bigl{[}(x_{1}+x_{2}+x_{3})^{3}x_{4}x_{5}^{3}\bigr{]}(\xi_{(1,2,3)}+\xi_{(1,3,2)}),\qquad\bigl{[}(x_{1}+x_{2}+x_{3})x_{4}^{3}x_{5}^{3}\bigr{]}(\xi_{(1,2,3)}+\xi_{(1,3,2)}),$
$\bigl{[}(x_{1}+x_{2}+x_{3})^{3}x_{4}^{2}x_{5}^{2}\bigr{]}(\xi_{(1,2,3)}+\xi_{(1,3,2)}),\qquad\bigl{[}(x_{1}+x_{2}+x_{3})^{2}x_{4}^{3}x_{5}^{2}\bigr{]}(\xi_{(1,2,3)}+\xi_{(1,3,2)}),$
$\bigl{[}(x_{1}+x_{2}+x_{3})^{2}x_{4}^{2}x_{5}^{3}\bigr{]}(\xi_{(1,2,3)}+\xi_{(1,3,2)})$
@@
This gives the following diamond:
[…]
Acknowledgements